OpenStudy (anonymous):

lim X-->infinity (e^4x +cosx -1) /2+e^x

OpenStudy (freckles):

I have you tried using squeeze theorem

OpenStudy (freckles):

we know -1<=cos(x)<=1 adding -1 on all sides gives -2<=cos(x)-1<=0 adding e^(4x) on all sides gives e^(4x)-2<=e^(4x)+cos(x)-1<=e^(4x) then of course it will not change the direction of the inequality if you divide by a positive function e^x

OpenStudy (anonymous):

so the answer is a range?

OpenStudy (freckles):

I didn't say that

OpenStudy (freckles):

I said to use squeeze theorem

OpenStudy (freckles):

that is why I set up an inequality

OpenStudy (freckles):

did you change your expression

OpenStudy (freckles):

earlier I seen \[\lim_{x \rightarrow \infty}\frac{e^{4x}+\cos(x)-1}{e^{x}}\]

OpenStudy (anonymous):

yes it was plus 2 at the bottome

OpenStudy (anonymous):

i though it didnt matter

OpenStudy (anonymous):

since after lhopital rule the 2 is gone

OpenStudy (freckles):

well e^x+2 is still positive so divide all side of the inequality I have up there by e^x+2 instead of e^x

OpenStudy (anonymous):

e^(4x)-2<=e^(4x)+cos(x)-1<=e^(4x) , all divided by (e^x +2)?

OpenStudy (freckles):

\[\frac{e^{4x}-2}{e^{x}+2} \le \frac{e^{4x}+\cos(x)-1}{e^{x}+2} \le \frac{e^{4x}}{e^x+2}\] yep now apply squeeze theorem here

OpenStudy (freckles):

also called sandwich theorem or pinching theorem

OpenStudy (anonymous):

infinity between infinity

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{e^{4x}}{e^x+2}=? \\ \lim_{x \rightarrow \infty}\frac{e^{4x}-2}{e^{4x}+2}=?\]

OpenStudy (freckles):

it is easier to find those limits

OpenStudy (freckles):

if those are the same then you determine the limit of the middle function

OpenStudy (freckles):

and yes that one is a type

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{e^{4x}}{e^x+2}=? \\ \lim_{x \rightarrow \infty}\frac{e^{4x}-2}{e^{x}+2}=? \]

OpenStudy (anonymous):

i got infinity - infinity form for the limits

OpenStudy (anonymous):

i see the cancellation of e^2x -2 , e^2x +2

OpenStudy (freckles):

since both have the same limit then the middle function also goes to infinity as x goes to infinity

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

e^4x is much bigger than e^x so they both go to infinity?